OpenStudy (anonymous):

Prove: cos θ - cos θ•sin2 θ = cos3 θ. You must show all work.

OpenStudy (anonymous):

\[\cos \theta-\cos \theta \times \sin^2 \theta=\cos^3 \theta\]

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ganeshie8

zepdrix (zepdrix):

Oh it's an exponent 3. I kept thinking it was a (3theta) lol

OpenStudy (anonymous):

That would make life complicated XD

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\cos \theta}-\color{orangered}{\cos \theta} \sin ^2\theta\quad=\quad\cos^3\theta\]Let's factor a cos(theta) out of each term.

zepdrix (zepdrix):

Get's us here, yes? \[\Large\rm \color{orangered}{\cos \theta}\left(1-\sin ^2\theta\right)\quad=\quad\cos^3\theta\]

OpenStudy (anonymous):

Ok, I'm with you

zepdrix (zepdrix):

Recall your Pythagorean Identity for sine and cosine:\[\Large\rm \sin^2\theta+\cos^2\theta=1\]If we subtract sin^2(theta) from each side, we get,\[\Large\rm \color{royalblue}{\cos^2\theta=1-\sin^2\theta}\] Hmmmmmm, can we apply that to our problem somehow?\[\Large\rm \cos \theta\left(\color{royalblue}{1-\sin ^2\theta}\right)\quad=\quad\cos^3\theta\]

OpenStudy (anonymous):

replace the 1 with sin^2+cos^2, then we can cancel out the sin^2, and when we distribute, we'll have cos^3

OpenStudy (anonymous):

Yayy!

OpenStudy (anonymous):

I think

zepdrix (zepdrix):

\[\Large\rm \cos \theta\left(\color{royalblue}{\cos^2\theta}\right)\quad=\quad\cos^3\theta\]

zepdrix (zepdrix):

Sorry I ran off for a sec c: Yayyy good job!

OpenStudy (anonymous):

Haha it's totally fine! Could you maybe help me with one more? I think it's relatively short

zepdrix (zepdrix):

Sure.

OpenStudy (anonymous):

f(x) = 3x^2 + 12x + 16 g(x) = 2 sin(2x - π) + 4 Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

If it were me, I'd just graph them both using tech like geogebra and look for the minimum. In and out in a minute. But that isn't what they want from me

zepdrix (zepdrix):

\[\Large\rm f(x)=3x^2+12x+16 \]The leading term (largest term) is a square. Second power tells us that this will be a parabola. I guess we would want to put this into vertex form in order to see the vertex (lowest point) easier.

OpenStudy (anonymous):

vertex form issssssss?

zepdrix (zepdrix):

\[\Large\rm f(x)=a(x-h)^2+k\]See how the x is written in terms of a perfect square? The vertex (lowest point) will be located at (h,k). So the minimum value will be k.

zepdrix (zepdrix):

\[\Large\rm f(x)=3x^2+12x+16\]Factor a 3 out of the first two terms,\[\Large\rm f(x)=3(\color{green}{x^2+4x})+16\]Now our goal is to try and complete the square on these green terms. When you have, \(\Large\rm x^2+bx\) To complete the square, You take half of your b coefficient, and square it. \(\Large\rm \left(\frac{b}{2}\right)^2\) That is the value that will complete the square.

OpenStudy (anonymous):

That's still 4

zepdrix (zepdrix):

So in our case, the b coefficient is 4. Half of 4 is 2, 2 squared is 4.\[\Large\rm f(x)=3(\color{green}{x^2+4x+4})+16\]So 4 is the value that will complete the square for us. But we can't just add 4 willy nilly like this. We have to also subtract 4 to keep the equation balanced.\[\Large\rm f(x)=3(\color{green}{x^2+4x+4}-4)+16\]We don't want this black 4 inside the parenthesis though. It isn't part of our perfect square. So we'll factor it out, multiplying by the 3.

zepdrix (zepdrix):

So it comes out like this, 3 times -4 \[\Large\rm f(x)=3(\color{green}{x^2+4x+4})-12+16\]

OpenStudy (anonymous):

Does that mean the minimum is 4?

zepdrix (zepdrix):

Yes very good! We can write our x stuff as a perfect square now, but that doesn't really matter.\[\Large\rm f(x)=3(\color{green}{x+2})^2-12+16\]\[\Large\rm f(x)=3(\color{green}{x+2})^2+4\]

OpenStudy (anonymous):

Thank you!!

zepdrix (zepdrix):

Remember that \(\Large\rm y=x^2\) is a parabola opening upwards. So we have the same thing going on with our f(x) except the vertex point has been shifted 2 to the right, and 4 up.

OpenStudy (anonymous):

ok :)

zepdrix (zepdrix):

\[\Large\rm g(x)=2\sin(2x-π)+4\]So this one... hmm let's see. So our general form: \[\Large\rm g(x)=A \sin\left[B(x-C)\right]+D\] Where the function will have an amplitude of \(\Large\rm A\), A period of \(\Large\rm \frac{2\pi}{B}\), is shifted horizontally by \(\Large\rm C\), and is shifted vertically by \(\Large\rm D\).

zepdrix (zepdrix):

Let's ignore the period and horizontal shift for now. Those won't affect the height of the function at all.

OpenStudy (anonymous):

So the minimum is 4?

OpenStudy (anonymous):

Or is that the Max?

zepdrix (zepdrix):

Here is our normal sine curve. It goes up to 1, down to negative 1.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nice graph by the way

zepdrix (zepdrix):

With an amplitude of 2, it will instead go up to 2 and down to -2.